# Stability of ranks under field extensions

### v3

@@ 5. Partition rank and slice rank under field extension
Suppose that $\mathbb{K}$ is either a finite field or a perfect infinite field and $\mathbb{F}$ is a finite extension of $\mathbb{K}$. Given any tensor $S\in\mathbb{F}^{n_{1}}\otimes\cdots\otimes\mathbb{F}^{n_{d}}$ with $\operatorname{PR}_{\mathbb{F}}(S)=r$, we have
$r\leq\operatorname{PR}_{\mathbb{K}}(S)\leq rR_{\mathbb{K}}(T_{3,\mathbb{F}}).$
@@
The first inequality follows immediately from the definition. Since $\operatorname{PR}_{\mathbb{F}}(S)=r$, we may write $S=\sum_{i=1}^{r}S_{i}$ for some $S_{i}\in\mathbb{F}^{n_{1}}\otimes\cdots\otimes\mathbb{F}^{n_{d}}$ such that $\operatorname{PR}_{\mathbb{F}}(S_{i})=1$. This implies that $\operatorname{PR}_{\mathbb{K}}(S)\leq\sum_{i=1}^{r}\operatorname{PR}_{\mathbb{K}}(S_{i})$ and it reduces to prove the lemma for $r=1$. If $\operatorname{PR}_{\mathbb{F}}(S)=1$, we may write $S=a\otimes b$ for some $a\in\mathbb{F}^{n_{\pi(1)}}\otimes\cdots\otimes\mathbb{F}^{n_{\pi(p)}}$ and $b\in\mathbb{F}^{n_{\pi(p+1)}}\otimes\cdots\otimes\mathbb{F}^{n_{\pi(d)}}$. Here $\pi\in\mathfrak{S}_{d}$ and $a\otimes b$ is identified with the tensor in $\mathbb{F}^{n_{1}}\otimes\cdots\otimes\mathbb{F}^{n_{d}}$ via the natural isomorphism
$\mathbb{F}^{n_{\pi(1)}}\otimes\cdots\otimes\mathbb{F}^{n_{\pi(d)}}\simeq\mathbb{F}^{n_{1}}\otimes\cdots\otimes\mathbb{F}^{n_{d}}.$
For convenience, we always adopt this convention in the sequel.
@@
We observe that $\mathbb{F}^{n_{\pi(1)}}\otimes\cdots\otimes\mathbb{F}^{n_{\pi(p)}}=\left(\mathbb{K}^{n_{\pi(1)}}\otimes\cdots\otimes\mathbb{K}^{n_{\pi(p)}}\right)\otimes_{\mathbb{K}}\mathbb{F}$, from which we may identify $a$ with a $\mathbb{K}$-multilinear map $f_{a}:\mathbb{K}^{n_{\pi(1)}}\times\cdots\times\mathbb{K}^{n_{\pi(p)}}\to\mathbb{F}$. Similarly, we identify $b$ (resp. $a\otimes b$) with a $\mathbb{K}$-multilinear map $f_{b}:\mathbb{K}^{n_{\pi(p+1)}}\times\cdots\times\mathbb{K}^{n_{\pi(d)}}\to\mathbb{F}$ (resp. $f_{a\otimes b}:\mathbb{K}^{n_{1}}\times\cdots\times\mathbb{K}^{n_{d}}\to\mathbb{F}$). We notice that as $K$-multilinear maps, it holds that $f_{a\otimes b}=T_{3,\mathbb{F}}\circ(f_{a}\otimes f_{b})$. Suppose that $T_{3,\mathbb{F}}=\sum_{j=1}^{s}\alpha_{j}\otimes\beta_{j}\otimes u_{j}$ where $s=R_{\mathbb{K}}(T_{3,\mathbb{F}})$, $\alpha_{j}$ and $\beta_{j}$ are $\mathbb{K}$-linear forms for $1\leq j\leq s$. Then we can write
@@
This implies that $\operatorname{PR}_{\mathbb{K}}(a\otimes b)\leq s=R_{\mathbb{K}}(T_{3,\mathbb{F}})$. ∎
@@
Proofs for slice rank and cp-rank are the same as that for partition rank, thus we only prove for partition rank. Denote $r\coloneqq\operatorname{PR}_{\overline{\mathbb{K}}}\left(T^{\overline{\mathbb{K}}}\right)$. Then for each $1\leq j\leq r$, there are $\pi\coloneqq\pi_{j}\in\mathfrak{S}_{d}$, $1\leq p\coloneqq p_{j}\leq d$, $a_{j}\in\overline{\mathbb{K}}^{n_{\pi(1)}}\otimes\cdots\otimes\overline{\mathbb{K}}^{n_{\pi(p)}}$ and $b_{j}\in\overline{\mathbb{K}}^{n_{\pi(p+1)}}\otimes\cdots\otimes\overline{\mathbb{K}}^{n_{\pi(d)}}$ such that
@@
$a_{j}\in\mathbb{F}^{n_{\pi(1)}}\otimes\cdots\otimes\mathbb{F}^{n_{\pi(p)}},\quad b_{j}\in\mathbb{F}^{n_{\pi(p+1)}}\otimes\cdots\otimes\mathbb{F}^{n_{\pi(d)}}.$ (19)
@@
$r=\operatorname{PR}_{\overline{\mathbb{K}}}\left(T^{\overline{\mathbb{K}}}\right)=\operatorname{PR}_{\mathbb{F}}\left(T^{\mathbb{F}}\right),\quad T^{\overline{\mathbb{K}}}=T^{\mathbb{F}}=T\boxtimes_{\mathbb{K}}T_{d,\mathbb{F}}.$
By Lemma 5.1, Proposition 3.1 and Theorem 3.2, there exists $c_{1}\coloneqq c_{1}(d,\mathbb{K})$ such that $\operatorname{PR}_{\mathbb{K}}(T\boxtimes_{\mathbb{K}}T_{d,\mathbb{F}})\leq c_{1}lr$. On the other side, Proposition 3.1 and Theorem 3.7 implies that there exists some constant $c_{2}\coloneqq c_{2}(d,\mathbb{K})$ such that $Q_{\mathbb{K}}(T_{d,\mathbb{F}})\geq c_{2}l$. This implies $T\boxtimes_{\mathbb{K}}T_{d,\mathbb{F}}\succeq T^{\oplus c_{2}l}$ and by the assumption we have
$c_{1}lr\geq\operatorname{PR}_{\mathbb{K}}(T\boxtimes_{\mathbb{K}}T_{d,\mathbb{F}})\geq cc_{2}l\operatorname{PR}_{\mathbb{K}}(T).$
@@
Let $T\in\mathbb{K}^{n_{1}}\otimes\cdots\otimes\mathbb{K}^{n_{d}}$ be a tensor. If $\mathbb{K}$ is a perfect infinite field, then we have
@@
[…]
where $\mathbb{K}[\mathsf{V}]$ denotes the ring of polynomials on a vector space $\mathsf{V}$. The image of this injection consists of all multilinear polynomials on $(\mathbb{K}^{m_{1}})^{\ast}\oplus\cdots\oplus(\mathbb{K}^{m_{s}})^{\ast}$. The following inequality is stated in [28, Theorem 5] with a sketched proof involving several lemmas. We provide below a shorter proof.
@@
Let $\mathbb{K}$ be a field and let $T\in\mathbb{K}^{n_{1}}\otimes\cdots\otimes\mathbb{K}^{n_{d}}$ be a tensor. Then $\operatorname{GR}_{\mathbb{K}}(T)\leq\operatorname{PR}_{\mathbb{K}}(T)$.
@@
Since $\operatorname{GR}_{\mathbb{K}}(T)=\operatorname{GR}_{\overline{\mathbb{K}}}\left(T^{\overline{\mathbb{K}}}\right)$ and $\operatorname{PR}_{\overline{\mathbb{K}}}\left(T^{\overline{\mathbb{K}}}\right)\leq\operatorname{PR}_{\mathbb{K}}(T)$, we may assume that $\mathbb{K}$ is algebraically closed. We define an $n_{d}$-dimensional subspace $\mathsf{W}\coloneqq\langle T,(\mathbb{K}^{n_{d}})^{\ast}\rangle$ in $\mathbb{K}^{n_{1}}\otimes\cdots\otimes\mathbb{K}^{n_{d-1}}$. Denote $\mathsf{V}\coloneqq(\mathbb{K}^{n_{1}})^{\ast}\oplus\cdots\oplus(\mathbb{K}^{n_{d-1}})^{\ast}$. Let $\tau$ be the injection defined in (21) and let $J$ be the ideal generated by $\tau(\mathsf{W})$ in $\mathbb{K}[\mathsf{V}]$. Obviously, the variety defined by $J$ coincides with $Z_{d}(T)$ defined in (4) and $\operatorname{GR}_{\mathbb{K}}(T)=\operatorname{codim}J$.
@@
Since $\mathsf{W}=\langle T,(\mathbb{K}^{n_{d}})^{\ast}\rangle=\sum_{i=1}^{r}a_{i}\otimes\langle b_{i},(\mathbb{K}^{n_{d}})^{\ast}\rangle$ and $J$ is generated by $\mathsf{W}$, we may derive $J\subseteq(f_{1},\dots,f_{r})$. According to [15, Theorem 10.2], we conclude that $\operatorname{codim}J\leq r$. ∎
@@
    for any $T\in\mathbb{F}_{q}^{n_{1}}\otimes\cdots\otimes\mathbb{F}_{q}^{n_{d}}$.
@@
    for any $T\in\mathbb{F}_{q}^{n_{1}}\otimes\cdots\otimes\mathbb{F}_{q}^{n_{d}}$.
@@
    for any $T\in\mathbb{F}_{q}^{n_{1}}\otimes\cdots\otimes\mathbb{F}_{q}^{n_{d}}$.
@@
    for any $T\in\mathbb{F}_{q}^{n_{1}}\otimes\cdots\otimes\mathbb{F}_{q}^{n_{d}}$.
@@
Thus, we have $\operatorname{PR}_{\mathbb{F}_{q}}\left(T^{\oplus n}\right)\geq\operatorname{GR}_{\mathbb{F}_{q}}\left(T^{\oplus n}\right)=n\operatorname{GR}_{\mathbb{F}_{q}}(T)$ by Lemma 5.5. This proves (b) $\implies$ (c). If (c) holds, then (d) also holds by Lemma 5.2. The implication (d) $\implies$ (a) is obtained by combining [9, Corollary 3] and the proof of [1, Theorem 1.13].
One direction of the ’moreover’ part is trivial. To prove the other direction, it is sufficient to prove that (c) holds for $q=p^{l}$ for any $l\in\mathbb{N}$ as long as it holds for $p$. Given a tensor $T\in\mathbb{F}_{q}^{n_{1}}\otimes\cdots\otimes\mathbb{F}_{q}^{n_{d}}$, Theorem 3.2 and Lemma 5.1 indicate the existence of a constant $C(d,p)$ such that $\operatorname{PR}_{\mathbb{F}_{p}}(T^{\oplus n})\leq C(d,p)l\operatorname{PR}_{\mathbb{F}_{q}}(T^{\oplus n})$. Since (c) holds for $p$, we also have $c(d,p)n\operatorname{PR}_{\mathbb{F}_{p}}(T)\leq\operatorname{PR}_{\mathbb{F}_{p}}(T^{\oplus n})$. We obtain

### v4

@@ 5 Partition rank and slice rank under field extension
Suppose that $\mathbb{K}$ is either a finite field or a perfect infinite field and $\mathbb{F}$ is a finite extension of $\mathbb{K}$. Given any $S\in\mathbb{F}^{n_{1}}\otimes_{\mathbb{F}}\cdots\otimes_{\mathbb{F}}\mathbb{F}^{n_{d}}$, we have
$\operatorname{PR}_{\mathbb{F}}(S)\leq\operatorname{PR}_{\mathbb{K}}(S_{\mathbb{K}})\leq\operatorname{R}_{\mathbb{K}}(M_{3,\mathbb{F}})\operatorname{PR}_{\mathbb{F}}(S).$
The same is true if we replace the partition rank in the above statement by slice rank or cp-rank, with the only modification that $\operatorname{R}_{\mathbb{K}}(M_{3,\mathbb{F}})$ is substituted by $\operatorname{R}_{\mathbb{K}}(M_{d+1,\mathbb{F}})$ for cp-rank.
@@
We first establish $\operatorname{PR}_{\mathbb{F}}(S)\leq\operatorname{PR}_{\mathbb{K}}(S_{\mathbb{K}})$. Since $\mathbb{F}$ is a separable extension of $\mathbb{K}$, we may write $\mathbb{F}=\mathbb{K}(\alpha)$ for some $\alpha\in\mathbb{F}$. For each $1\leq k\leq d$, we let $\{x^{(k)}_{j}:1\leq j\leq n_{k}\}$ be an $\mathbb{F}$-basis of $\mathbb{F}^{n_{k}}$. Then $\{\alpha^{s}x^{(k)}_{j}:0\leq s\leq l-1,\;1\leq j\leq n_{k}\}$ is a $\mathbb{K}$-basis of $\mathbb{F}^{n_{k}}$, where $l\coloneqq[\mathbb{F}:\mathbb{K}]$. Thus, we may write
$S_{\mathbb{K}}=\sum_{\begin{subarray}{c}j_{1},\dots,j_{d}\\ s_{1},\dots,s_{d}\end{subarray}}S^{j_{1},\dots,j_{d}}_{s_{1},\dots,s_{d}}\left(\alpha^{s_{1}}x^{(1)}_{j_{1}}\right)\otimes_{\mathbb{K}}\cdots\otimes_{\mathbb{K}}\left(\alpha^{s_{d}}x^{(d)}_{j_{d}}\right)\in\mathbb{F}^{n_{1}}\otimes_{\mathbb{K}}\cdots\otimes_{\mathbb{K}}\mathbb{F}^{n_{d}}.$
where $S^{j_{1},\dots,j_{d}}_{s_{1},\dots,s_{d}}\in\mathbb{K}$, $1\leq j_{k}\leq n_{k}$ and $0\leq s_{k}\leq l-1$ for each $1\leq k\leq d$. By definition of $S_{\mathbb{K}}$, we have
$S((x^{(1)}_{j_{1}})^{\ast},\dots,(x^{(d-1)}_{j_{d-1}})^{\ast})=S_{\mathbb{K}}(x^{(1)}_{j_{1}},\dots,x^{(d-1)}_{j_{d-1}})=\sum_{s=0,j_{d}=1}^{l-1,n_{d}}S^{j_{1},\dots,j_{d}}_{0,\dots,0,s}\alpha^{s}x^{(d)}_{j_{d}}$.
Here $\{(x^{(k)}_{j})^{\ast}:1\leq j\leq n_{k}\}$ is the $\mathbb{F}$-basis of $(\mathbb{F}^{n_{k}})^{\ast}$ dual to $\{x^{(k)}_{j}:1\leq j\leq n_{k}\}$ for $1\leq k\leq d-1$. Suppose $\{y^{(k)}_{s_{k},j_{k}}:0\leq s_{k}\leq l-1,\;1\leq j\leq n_{k}\}$ is an $\mathbb{F}$-basis of $\mathbb{F}^{ln_{k}}$ with the dual basis $\{(y^{(k)}_{s_{k},j_{k}})^{\ast}:0\leq s_{k}\leq l-1,\;1\leq j\leq n_{k}\}$. We consider an $\mathbb{F}$-tensor:
$(S_{\mathbb{K}})^{\mathbb{F}}=\sum_{\begin{subarray}{c}j_{1},\dots,j_{d}\\ s_{1},\dots,s_{d}\end{subarray}}S^{j_{1},\dots,j_{d}}_{s_{1},\dots,s_{d}}y^{(1)}_{s_{1},j_{1}}\otimes_{\mathbb{F}}\cdots\otimes_{\mathbb{F}}y^{(d)}_{s_{d},j_{d}}\in\mathbb{F}^{ln_{1}}\otimes_{\mathbb{F}}\cdots\otimes_{\mathbb{F}}\mathbb{F}^{ln_{d}}.$
For $1\leq k\leq d-1$, we let $\tau_{k}:\left(\mathbb{F}^{n_{k}}\right)^{\ast}\to\left(\mathbb{F}^{ln_{k}}\right)^{\ast}$ be the $\mathbb{F}$-linear map obtained by linearly extending the assignment $(x^{(k)}_{j})^{\ast}\mapsto(y^{(k)}_{0,j})^{\ast}$, $1\leq j\leq n_{k}$. We also define $p:\mathbb{F}^{ln_{d}}\to\mathbb{F}^{n_{d}}$ to be the $\mathbb{F}$-linear map obtained by linearly extending the assignment $y^{(d)}_{s,j}\mapsto\alpha^{s}x^{(d)}_{j}$, $1\leq j\leq n_{d}$, $0\leq s\leq l-1$. We claim the the following digram commutes:
$(\mathbb{F}^{n_{1}})^{\ast}\times\cdots\times(\mathbb{F}^{n_{d-1}})^{\ast}$$(\mathbb{F}^{ln_{1}})^{\ast}\times\cdots\times(\mathbb{F}^{ln_{d-1}})^{\ast}$$\mathbb{F}^{n_{d}}$$\mathbb{F}^{ln_{d}}$$\scriptstyle{\tau_{1}\times\cdots\times\tau_{d-1}}$$\scriptstyle{S}$$\scriptstyle{{S^{\mathbb{F}}}}$$\scriptstyle{p}$
Indeed, it is straightforward to verify that
$p\left((S_{\mathbb{K}})^{\mathbb{F}}\left(\tau_{1}((x^{(1)}_{j_{1}})^{\ast}),\dots,\tau_{d-1}((x^{(d-1)}_{j_{d-1}})^{\ast})\right)\right)$ $=p\left((S_{\mathbb{K}})^{\mathbb{F}}\left((y^{(1)}_{0,j_{1}})^{\ast},\dots,(y^{(d-1)}_{0,j_{d-1}})^{\ast}\right)\right)$
$=p\left(\sum_{s=0,j_{d}=1}^{l-1,n_{d}}S^{j_{1},\dots,j_{d}}_{0,\dots,0,s}y^{(d)}_{s,j_{d}}\right)$
$=\sum_{s=0,j_{d}=1}^{l-1,n_{d}}S^{j_{1},\dots,j_{d}}_{0,\dots,0,s}\alpha^{s}x^{(d)}_{j_{d}}$
$=S\left((x^{(1)}_{j_{1}})^{\ast},\dots,(x^{(d-1)}_{j_{d-1}})^{\ast}\right)$.
Thus, we may conclude that $\operatorname{PR}_{\mathbb{K}}(S_{\mathbb{K}})\geq\operatorname{PR}_{\mathbb{F}}((S_{\mathbb{K}})^{\mathbb{F}})\geq\operatorname{PR}_{\mathbb{F}}(S)$.
@@
Next, we prove $\operatorname{PR}_{\mathbb{K}}(S_{\mathbb{K}})\leq\operatorname{R}_{\mathbb{K}}(M_{3,\mathbb{F}})\operatorname{PR}_{\mathbb{F}}(S)$. Denote $r\coloneqq\operatorname{PR}_{\mathbb{F}}(S)$. Then we may write $S=\sum_{i=1}^{r}S_{i}$ for some $S_{i}\in\mathbb{F}^{n_{1}}\otimes_{\mathbb{F}}\cdots\otimes_{\mathbb{F}}\mathbb{F}^{n_{d}}$ such that $\operatorname{PR}_{\mathbb{F}}(S_{i})=1$. This implies that $\operatorname{PR}_{\mathbb{K}}(S)\leq\sum_{i=1}^{r}\operatorname{PR}_{\mathbb{K}}(S_{i})$ and it reduces to prove the inequality for $r=1$. If $\operatorname{PR}_{\mathbb{F}}(S)=1$, we may write $S=a\otimes b$ for some $a\in\mathbb{F}^{n_{\pi(1)}}\otimes_{\mathbb{F}}\cdots\otimes_{\mathbb{F}}\mathbb{F}^{n_{\pi(p)}}$ and $b\in\mathbb{F}^{n_{\pi(p+1)}}\otimes_{\mathbb{F}}\cdots\otimes_{\mathbb{F}}\mathbb{F}^{n_{\pi(d)}}$. Here $\pi\in\mathfrak{S}_{d}$ and $a\otimes b$ is identified with the tensor in $\mathbb{F}^{n_{1}}\otimes_{\mathbb{F}}\cdots\otimes_{\mathbb{F}}\mathbb{F}^{n_{d}}$ via the natural isomorphism
$\mathbb{F}^{n_{\pi(1)}}\otimes_{\mathbb{F}}\cdots\otimes_{\mathbb{F}}\mathbb{F}^{n_{\pi(d)}}\simeq\mathbb{F}^{n_{1}}\otimes_{\mathbb{F}}\cdots\otimes_{\mathbb{F}}\mathbb{F}^{n_{d}}.$
@@
For convenience, we always adopt this convention in the sequel. We observe that $\mathbb{F}^{n_{\pi(1)}}\otimes_{\mathbb{F}}\cdots\otimes_{\mathbb{F}}\mathbb{F}^{n_{\pi(p)}}=\left(\mathbb{K}^{n_{\pi(1)}}\otimes_{\mathbb{K}}\cdots\otimes_{\mathbb{K}}\mathbb{K}^{n_{\pi(p)}}\right)\otimes_{\mathbb{K}}\mathbb{F}$, from which we may identify $a$ with a $\mathbb{K}$-multilinear map $f_{a}:\mathbb{K}^{n_{\pi(1)}}\times\cdots\times\mathbb{K}^{n_{\pi(p)}}\to\mathbb{F}$. Similarly, we identify $b$ (resp. $a\otimes b$) with a $\mathbb{K}$-multilinear map $f_{b}:\mathbb{K}^{n_{\pi(p+1)}}\times\cdots\times\mathbb{K}^{n_{\pi(d)}}\to\mathbb{F}$ (resp. $f_{a\otimes b}:\mathbb{K}^{n_{1}}\times\cdots\times\mathbb{K}^{n_{d}}\to\mathbb{F}$). We notice that as $\mathbb{K}$-multilinear maps, it holds that $f_{a\otimes b}=M_{3,\mathbb{F}}\circ(f_{a}\otimes f_{b})$. Suppose that $M_{3,\mathbb{F}}=\sum_{j=1}^{s}\alpha_{j}\otimes\beta_{j}\otimes u_{j}$ where $s=R_{\mathbb{K}}(M_{3,\mathbb{F}})$, $\alpha_{j}$ and $\beta_{j}$ are $\mathbb{K}$-linear forms for $1\leq j\leq s$. Then we can write
@@
This implies that $\operatorname{PR}_{\mathbb{K}}(a\otimes b)\leq s=R_{\mathbb{K}}(M_{3,\mathbb{F}})$.
To prove the two inequalities for slice rank and cp-rank, one can simply repeat the above argument while replacing partition rank and the corresponding decomposition with those of slice rank and cp-rank, respectively. ∎
Remark 5.2.
It is proved in [13, Proof of Corollary 1] that
$\operatorname{PR}_{\mathbb{K}}(T)\leq l\operatorname{PR}_{\mathbb{F}}(T^{\mathbb{F}}),$ (19)
where $\mathbb{K}$ is a finite field, $\mathbb{F}$ is a degree $l$ extension of $\mathbb{K}$ and $T$ is a $\mathbb{K}$-tensor. It is worth noticing that (19) is different from the one established in Lemma 5.1, although they appear similar. The main difference is that (19) is concerned with the relation between partition ranks of a $\mathbb{K}$-tensor and its $\mathbb{F}$-extension, whereas Lemma 5.1 relates the partition rank of an $\mathbb{F}$-tensor and that of its $\mathbb{K}$-restriction.
On the other hand, one of the anonymous reviewers has pointed out that the inequality $\operatorname{PR}_{\mathbb{K}}(S_{\mathbb{K}})\leq l\operatorname{PR}_{\mathbb{F}}(S)$ holds. Indeed, according to Remark 4.1, the $\mathbb{K}$-multilinear function corresponding to $S_{\mathbb{K}}$ is actually ${\operatorname{tr}}\circ f_{a\otimes b}$, where we assume $S=a\otimes b$ as in the proof of Lemma 5.1. Since $\operatorname{rank}\left({\operatorname{tr}}\circ M_{3,\mathbb{F}}\right)=l$ and
${\operatorname{tr}}\circ f_{a\otimes b}=\left({\operatorname{tr}}\circ M_{3,\mathbb{F}}\right)\circ(f_{a}\otimes f_{b}),$
the desired inequality follows immediately. The same argument applies to slice rank and cp-rank. Consequently, we may improve the constant factor from $\operatorname{R_{\mathbb{K}}}(M_{d+1,\mathbb{F}})$ to $\operatorname{R_{\mathbb{K}}}(M_{d,\mathbb{F}})$.
@@
Proofs for slice rank and cp-rank are the same as that for partition rank, thus we only prove for partition rank. Denote $r\coloneqq\operatorname{PR}_{\overline{\mathbb{K}}}\left(T^{\overline{\mathbb{K}}}\right)$. Then for each $1\leq j\leq r$, there are $\pi\coloneqq\pi_{j}\in\mathfrak{S}_{d}$, $1\leq p\coloneqq p_{j}\leq d$, $a_{j}\in\overline{\mathbb{K}}^{n_{\pi(1)}}\otimes_{\overline{\mathbb{K}}}\cdots\otimes_{\overline{\mathbb{K}}}\overline{\mathbb{K}}^{n_{\pi(p)}}$ and $b_{j}\in\overline{\mathbb{K}}^{n_{\pi(p+1)}}\otimes_{\overline{\mathbb{K}}}\cdots\otimes_{\overline{\mathbb{K}}}\overline{\mathbb{K}}^{n_{\pi(d)}}$ such that
@@
$a_{j}\in\mathbb{F}^{n_{\pi(1)}}\otimes_{\mathbb{F}}\cdots\otimes_{\mathbb{F}}\mathbb{F}^{n_{\pi(p)}},\quad b_{j}\in\mathbb{F}^{n_{\pi(p+1)}}\otimes_{\mathbb{F}}\cdots\otimes_{\mathbb{F}}\mathbb{F}^{n_{\pi(d)}}.$ (21)
@@
$r=\operatorname{PR}_{\overline{\mathbb{K}}}\left(T^{\overline{\mathbb{K}}}\right)=\operatorname{PR}_{\mathbb{F}}\left(T^{\mathbb{F}}\right),\quad(T^{\mathbb{F}})_{\mathbb{K}}=T\boxtimes_{\mathbb{K}}M_{d,\mathbb{F}}.$
By Lemma 5.1, Proposition 3.1 and Theorem 3.2, there exists $c_{1}\coloneqq c_{1}(d,\mathbb{K})$ such that $\operatorname{PR}_{\mathbb{K}}(T\boxtimes_{\mathbb{K}}M_{d,\mathbb{F}})\leq c_{1}lr$. On the other side, Proposition 3.1 and Theorem 3.7 implies that there exists some constant $c_{2}\coloneqq c_{2}(d,\mathbb{K})$ such that $\operatorname{Q}_{\mathbb{K}}(M_{d,\mathbb{F}})\geq c_{2}l$. This implies $T\boxtimes_{\mathbb{K}}M_{d,\mathbb{F}}\succeq T^{\oplus c_{2}l}$ and by the assumption we have
$c_{1}lr\geq\operatorname{PR}_{\mathbb{K}}(T\boxtimes_{\mathbb{K}}M_{d,\mathbb{F}})\geq cc_{2}l\operatorname{PR}_{\mathbb{K}}(T).$
@@
Let $T\in\mathbb{K}^{n_{1}}\otimes_{\mathbb{K}}\cdots\otimes_{\mathbb{K}}\mathbb{K}^{n_{d}}$ be a tensor. If $\mathbb{K}$ is a perfect infinite field, then we have
@@
$\mathbb{K}^{m_{1}}\otimes_{\mathbb{K}}\cdots\otimes_{\mathbb{K}}\mathbb{K}^{m_{t}}$$\mathbb{K}\left[(\mathbb{K}^{m_{1}})^{\ast}\oplus\cdots\oplus(\mathbb{K}^{m_{s}})^{\ast}\right]$$\operatorname{Hom}\left((\mathbb{K}^{m_{1}})^{\ast}\otimes_{\mathbb{K}}\cdots\otimes_{\mathbb{K}}(\mathbb{K}^{m_{t}})^{\ast},\mathbb{K}\right)$$\mathbb{K}\left[(\mathbb{K}^{m_{1}})^{\ast}\oplus\cdots\oplus(\mathbb{K}^{m_{t}})^{\ast}\right]$$\scriptstyle{\tau}\scriptstyle{\simeq}$ (23)
where $\mathbb{K}[\mathsf{V}]$ denotes the ring of polynomials on a vector space $\mathsf{V}$. The image of this injection consists of all multilinear polynomials on $(\mathbb{K}^{m_{1}})^{\ast}\oplus\cdots\oplus(\mathbb{K}^{m_{s}})^{\ast}$. The following inequality is stated in [35, Theorem 5] with a sketched proof involving several lemmas. We provide below a shorter proof.
@@
Let $\mathbb{K}$ be a field and let $T\in\mathbb{K}^{n_{1}}\otimes_{\mathbb{K}}\cdots\otimes_{\mathbb{K}}\mathbb{K}^{n_{d}}$ be a tensor. Then $\operatorname{GR}_{\mathbb{K}}(T)\leq\operatorname{PR}_{\mathbb{K}}(T)$.
@@
Since $\operatorname{GR}_{\mathbb{K}}(T)=\operatorname{GR}_{\overline{\mathbb{K}}}\left(T^{\overline{\mathbb{K}}}\right)$ and $\operatorname{PR}_{\overline{\mathbb{K}}}\left(T^{\overline{\mathbb{K}}}\right)\leq\operatorname{PR}_{\mathbb{K}}(T)$, we may assume that $\mathbb{K}$ is algebraically closed. We define an $n_{d}$-dimensional subspace $\mathsf{W}\coloneqq\langle T,(\mathbb{K}^{n_{d}})^{\ast}\rangle$ in $\mathbb{K}^{n_{1}}\otimes_{\mathbb{K}}\cdots\otimes_{\mathbb{K}}\mathbb{K}^{n_{d-1}}$. Denote $\mathsf{V}\coloneqq(\mathbb{K}^{n_{1}})^{\ast}\oplus\cdots\oplus(\mathbb{K}^{n_{d-1}})^{\ast}$. Let $\tau$ be the injection defined in (23) and let $J$ be the ideal generated by $\tau(\mathsf{W})$ in $\mathbb{K}[\mathsf{V}]$. Obviously, the variety defined by $J$ coincides with $Z_{d}(T)$ defined in (4) and $\operatorname{GR}_{\mathbb{K}}(T)=\operatorname{codim}J$.
@@
Since $\mathsf{W}=\langle T,(\mathbb{K}^{n_{d}})^{\ast}\rangle=\sum_{i=1}^{r}a_{i}\otimes\langle b_{i},(\mathbb{K}^{n_{d}})^{\ast}\rangle$ and $J$ is generated by $\mathsf{W}$, we may derive $J\subseteq(f_{1},\dots,f_{r})$. According to [19, Theorem 10.2], we conclude that $\operatorname{codim}J\leq r$. ∎
@@
    for any $T\in\mathbb{F}_{q}^{n_{1}}\otimes_{\mathbb{F}_{q}}\cdots\otimes_{\mathbb{F}_{q}}\mathbb{F}_{q}^{n_{d}}$.
@@
    for any $T\in\mathbb{F}_{q}^{n_{1}}\otimes_{\mathbb{F}_{q}}\cdots\otimes_{\mathbb{F}_{q}}\mathbb{F}_{q}^{n_{d}}$.
@@
    for any $T\in\mathbb{F}_{q}^{n_{1}}\otimes_{\mathbb{F}_{q}}\cdots\otimes_{\mathbb{F}_{q}}\mathbb{F}_{q}^{n_{d}}$.
@@
    for any $T\in\mathbb{F}_{q}^{n_{1}}\otimes_{\mathbb{F}_{q}}\cdots\otimes_{\mathbb{F}_{q}}\mathbb{F}_{q}^{n_{d}}$.
@@
Thus, we have $\operatorname{PR}_{\mathbb{F}_{q}}\left(T^{\oplus n}\right)\geq\operatorname{GR}_{\mathbb{F}_{q}}\left(T^{\oplus n}\right)=n\operatorname{GR}_{\mathbb{F}_{q}}(T)$ by Lemma 5.6. This proves (b) $\implies$ (c). If (c) holds, then (d) also holds by Lemma 5.3. The implication (d) $\implies$ (a) is obtained by combining [13, Corollary 3] and the proof of [1, Theorem 1.13].
One direction of the ‘moreover’ part is trivial. To prove the other direction, it is sufficient to prove that (c) holds for $q=p^{l}$ for any $l\in\mathbb{N}$ as long as it holds for $p$. Given a tensor $T\in\mathbb{F}_{q}^{n_{1}}\otimes_{\mathbb{F}_{q}}\cdots\otimes_{\mathbb{F}_{q}}\mathbb{F}_{q}^{n_{d}}$, Theorem 3.2 and Lemma 5.1 indicate the existence of a constant $C(d,p)$ such that $\operatorname{PR}_{\mathbb{F}_{p}}(T^{\oplus n})\leq C(d,p)l\operatorname{PR}_{\mathbb{F}_{q}}(T^{\oplus n})$. Since (c) holds for $p$, we also have $c(d,p)n\operatorname{PR}_{\mathbb{F}_{p}}(T)\leq\operatorname{PR}_{\mathbb{F}_{p}}(T^{\oplus n})$. We obtain